# Judging the Judges: A Systematic Study of Position Bias in LLM-as-a-Judge

### v8

@@ D.2 DevBench
Figure 6: Baseline comparisons of position bias of LLM judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse $PC$ of other models compared to it. Figure (b) utilizes an absolute $PF$ baseline of 0 and depicts preference fairness performances of LLM judges across tasks. Similar to findings on MTBench, position bias is significantly judge-dependent and task-dependent on DevBench as well.
@@ E.2 Test results
We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias. Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively. Table 5, 6 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively. Table 7, 8 serves for DevBench, and Table 3, 4 is conducted on the integrated set of both benchmarks. The impact of variables on the model is ranked from highest to lowest, from bottom to top. Removed variables listed as None indicate the full model at this given step.
@@ F.1 Judges, Candidates, and Tasks
The model names are exactly what MTBench Zheng et al. (2024b) and DevBench Li et al. (2024a) used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
@@ F.2 Prompt Settings
  • Chain-of-Thought Wei et al. (2022) prompts requiring the LLM judge to provide reasons for its judgment

### v9

@@ Appendix D More Results of Position Bias and Answer Quality Gap Measurement
Table 3 presents additional evaluation results for more open-source models, reinforcing our earlier observations. The scores consistently show that position bias varies across judges and tasks, while high repetition stability confirms that such biases are systematic rather than random. These extended results further validate our evaluation findings.
Model MTBench Results DevBench Results
PCPC PFPF RSRS Error PCPC PFPF RSRS Error
GPT-OSS-20B 0.79 $\pm$ 0.16 -0.04 $\pm$ 0.13 0.89 $\pm$ 0.17 0.06 0.84 $\pm$ 0.16 -0.04 $\pm$ 0.14 0.81 $\pm$ 0.25 0.08
Gemma-3n-E4B 0.68 $\pm$ 0.17 -0.07 $\pm$ 0.18 0.90 $\pm$ 0.17 0.00 0.41 $\pm$ 0.24 0.40 $\pm$ 0.35 0.93 $\pm$ 0.12 0.08
Llama-4-Scout-17B 0.79 $\pm$ 0.16 0.05 $\pm$ 0.15 0.95 $\pm$ 0.08 0.00 0.12 $\pm$ 0.18 0.88 $\pm$ 0.18 0.98 $\pm$ 0.03 0.00
Mistral-Small-24B 0.72 $\pm$ 0.19 -0.02 $\pm$ 0.17 0.87 $\pm$ 0.20 0.05 0.57 $\pm$ 0.18 0.27 $\pm$ 0.19 0.82 $\pm$ 0.31 0.02
Qwen2.5-72B 0.77 $\pm$ 0.17 -0.01 $\pm$ 0.16 0.94 $\pm$ 0.09 0.00 0.81 $\pm$ 0.15 0.11 $\pm$ 0.16 0.93 $\pm$ 0.13 0.00
Qwen2.5-Coder-32B 0.74 $\pm$ 0.18 0.01 $\pm$ 0.17 0.91 $\pm$ 0.14 0.00 0.81 $\pm$ 0.16 0.09 $\pm$ 0.17 0.92 $\pm$ 0.13 0.03
Qwen2.5-7B 0.55 $\pm$ 0.18 -0.21 $\pm$ 0.23 0.87 $\pm$ 0.22 0.00 0.54 $\pm$ 0.21 -0.45 $\pm$ 0.21 0.93 $\pm$ 0.12 0.00
Table 3: More evaluation results for Position Consistency (PCPC), Preference Fairness (PFPF), and Repetition Stability ($RS$) across MTBench and DevBench datasets.
@@ D.2 DevBench
Figure 6: Baseline comparisons of position bias of LLM judges across tasks on DevBench. An asterisk marks the statistical significance by Student’s t-tests. Figure (a) selects GPT-4 as the baseline, where asterisk demonstrates signficantly better or worse PCPC of other models compared to it. Figure (b) utilizes an absolute PFPF baseline of 0 and depicts preference fairness performances of LLM judges across tasks. Similar to findings on MTBench, position bias is significantly judge-dependent and task-dependent on DevBench as well.
@@ E.2 Test results
We operated bidirectional stepwise regression on both benchmarks individually and together to identify the factors that are significantly contributing to position bias. Specifically, the variables include lengths (input, output, and prompt), answer quality gap, LLM judges, candidate models, and task categories to predict position consistency and preference fairness respectively. Table 6, 7 records the results of final step in stepwise regression for predicting $PC$ and $PF$, respectively. Table 8, 9 serves for DevBench, and Table 4, 5 is conducted on the integrated set of both benchmarks. The impact of variables on the model is ranked from highest to lowest, from bottom to top. Removed variables listed as None indicate the full model at this given step.
@@ F.1 Judges, Candidates, and Tasks
The model names are exactly what MTBench zheng2024judging and DevBench li2024devbench used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
@@ F.2 Prompt Settings
  • Chain-of-Thought wei2022chain prompts requiring the LLM judge to provide reasons for its judgment